OpenStudy (anonymous):

One solution of the equation 6x^2-19x+3=0 is 1/6 What is the other solution?

OpenStudy (anonymous):

Please help!!

OpenStudy (anonymous):

Let f(x)=6x^2-19x+3 If f(1/6)=0, then (6x-1) is a factor of f(x) f(x)=(6x-1)(Ax-B) =6Ax^2-Ax-6Bx+B Comparing coefficients 6A=6 -A-6B=19 B=3 Just solve A and B, Ax-B=0 x=B/A